OpenStudy (anonymous):

Solve 3x = –24. A. –21 B. –12 C. –27 D. –8

OpenStudy (anonymous):

A @pooja195

pooja195 (pooja195):

-21/3=?

pooja195 (pooja195):

A is incorrect we are trying to isolate x in this case we divide both sides by 3 3/3=1 so x -21/3=?

OpenStudy (anonymous):

D

pooja195 (pooja195):

ooppsit wasnt 21 lol sorry but yes thats correct

pooja195 (pooja195):

I haye touchscreens -,-

OpenStudy (anonymous):

lol